CrCl3(s) + AgNO3 (aq)--> AgCl(s) + Cr(NO3)3(aq)
^^ unbalanced.

A 0.810 g sample of impure CrCl3 was dissolved in water and required 37.0 mL of 0.125 M AgNO3 solution to react with the CrCl3 in the sample. What was the mass percent of CrCl3 in the sample?

I think the balances version would be
CrCl3(s) + 3AgNO3 (aq)--> 3AgCl(s) + Cr(NO3)3(aq)
but that's as far as I can get.
1
Expert's answer
2012-03-01T09:44:41-0500
in 1000 ml ----------0.125 moles
in 37.0----------------x

x = 0.004625 moles

x 0.004625
CrCl3(s) + 3AgNO3 (aq)--> 3AgCl(s) + Cr(NO3)3(aq)

x=0.001541 moles

m = n*MW = 0.001541*112.89 = 0.1894g

w = 0.1894/0.810 * 100% = 23.38%